A cube of wood supporting a 200gm mass just float in water.when the mass is removed the cube rises by 2cm at equilibrium. Find side of the cube

Grade:11

1 Answers

Arun
6 years ago
Dear student
 
Please find the answer to your question
Let the edge of cube be l. When mass is on the cube of wood
200 g + l3 d wood g = l3 dwater 
⇒ l3 dwood = l3 dwater – 200 … (i)
When the mass is removed 
l3 dwood = (l - 2 ) l2 dwater … (ii)
From (i) and (ii) 
l3 dwater - 200 = (l- 2) l2 dwater
But dwater = 1 
∴ l3 - 200 = l2 (l - 2)
⇒ l = 10 cm
Regards
Arun (askIITians forum expert)